9/5X+ 32 = (X-32) 5/9
Solve for X. It is the temperature that cellius and Fahrenheit Are equal.
Show steps.

Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

9/5X+ 32 = (X-32) 5/9
Solve for X. It is the temperature that cellius [sic] and Fahrenheit Are equal.
-----------
9x/5 + 32 = (X-32)*5/9
Multiply by 45
81x + 1440 = 25*(x-32) = 25x - 800
56x = -2240
x = -40
-40 C = -40 F
===============
C = Celsius, not cellius
=============================
Do it like this:
F = 9C/5 + 32
@ F = C:
F = 9F/5 + 32
5F = 9F + 160
-4F = 160
F = -40 = C
-------
At -40, mercury freezes. -38.83 C, actually.

Answer by rothauserc(4718)

(9/5) * x +32 = (x-32) * (5/9)
:
9x/5 +32 = 5x/9 -160/9
:
45 is the least common multiple, multiply both sides of = by 45
:
81x +1440 = 25x -800
:
56x = -2240
:
***************************************
x = -40
:
-40 degrees Fahrenheit = -40 Centigrade
:
***************************************
: